# Low-Rank Tensor Decompositions for the Theory of Neural Networks

### v1

@@ I-1 NNs with low-rank weights
Many works represent weights in different NN architectures as low-rank tensors for their compression [8], and also for the compression of their gradient updates (such as in LoRA [9]). We will present results on the sensitivity [10] and convergence [11] of NN weight compression algorithms, on the generalization performance of such models [12, 13], and on their implicit biases. This will be covered in Section IV.
@@ I-3 Learning with differentiation
Through the connection between NNs and tensors obtained by derivatives, the theory of low-rank decompositions helps to establish polynomial time training algorithms, generalization performance guarantees, handling trainable activation functions, and parameter identifiability for NNs. We will discuss results for learning 2- and 3-layer NNs that use this connection by leveraging different low-rank matrix and tensor formats [20, 21, 22, 23, 24]. This will be covered in Section VI.
@@ II-D1 Generic properties
A wealth of results exist on generic properties of tensor decompositions, that is, the properties which hold for almost all tensors (as explained below) of a given size or tensor format. Generic properties can be studied using the powerful tools of algebraic geometry [4]. A key example is the generic tensor rank: all complex-valued tensors of given size $I_{1}\times\cdots\times I_{d}$ have the same CP rank $r_{\rm gen}\triangleq r_{\rm gen}(I_{1},\ldots,I_{d})$, except for a set of Lebesgue measure zero. In other words, if a complex tensor of a given size is drawn according to continuous probability distribution, then with probability one the tensor has CP rank $r_{\rm gen}$. For example, for cubic tensors $r_{\rm gen}(I,I,I)=\big{\lceil}\frac{I^{3}}{3I-2}\big{\rceil}$, and for many different tensor shapes the generic ranks are known. The generic ranks also give a bound $r_{\rm max}\leq 2r_{\rm gen}$ on maximal rank $r_{\rm max}$ of a given tensor size, which is otherwise difficult to study and unknown for most cases. For real-valued tensors, several typical ranks can occur (corresponding to sets of positive Lebesque measure), but the smallest typical rank is equal to $r_{\rm gen}$, see [4, 2, 18, 29] for an overview.
@@ II-D3 Other factorizations
Despite being widely used to explain the behavior of NNs, the CPD has limited flexibility, which is one reason many results only address the 2-layer case (see, e.g., [20, 34]). Different generalizations of the CPD exist: symmetric, partially symmetric (where symmetry is imposed on factors of the rank-one components), block-term decomposition (BTD, decomposition of a tensor as a sum of low-rank tensors); and the so-called tubal tensor rank decomposition, or t-SVD (often used for multichannel/3D data) which is a special case of the BTD with a fixed factor matrix [35]. The above decompositions are additive, and most of them fall under the umbrella of X-rank decompositions [18, 29]. Many generic properties of such decompositions (e.g., identifiability) can be studied using the tools from algebraic geometry [4, 29], similarly to the case of CPD. Besides additive decompositions, there are other generalizations of the CPD such as the Paratuck-2 [3], which enjoys uniqueness guarantees and is useful in the analysis of deep learning models [23].
@@ III Neural networks: basic setup
$\boldsymbol{h}_{\ell}$ $=f_{\ell}(\boldsymbol{h}_{\ell-1})$
$\triangleq\sigma_{\ell}(\boldsymbol{A}_{\ell}\boldsymbol{h}_{\ell-1}+\boldsymbol{b}_{\ell})\,,\qquad\text{(hidden layers)}$ (5)
@@
[…]
@@ IV-A Compression of weights
Note that tensorization has been used not only to compress NNs but also to increase their expressiveness, as a means of capturing higher order interactions between inputs/hidden variables. This was used in tensor attention modules in transformers [37] and in second-order RNNs [15], which were proven to have increased expressiveness over their original counterparts. However, the most common use of low-rank representations is still to reduce not only storage costs but also training and inference time. This is because in many cases the computations can be performed using the decomposed format itself, without need to reconstruct the full weight tensor (which would not decrease the computation time during inference).
The availability of different low-rank formats introduces the question of what is the best loss function or tensor rank for compressing a given NN architecture. Relevant work in this direction include [10], which proposed stable loss functions to avoid degenerate components when compressing CNN layers, and [11], which uses the Riemannian geometry of the Tucker decomposition to train NNs in compressed format with improved convergence performance, besides proposing an adaptive rank selection strategy.
@@ V-A2 Polynomial neural networks
$f(\boldsymbol{x})=\boldsymbol{A}_{2}\sigma_{1}(\boldsymbol{A}_{1}\boldsymbol{x})\,\iff\mathcal{F}\emph{ has a CPD with factors }\boldsymbol{A}_{2},\underbrace{\boldsymbol{A}_{1}^{\top},\ldots,\boldsymbol{A}^{\top}_{1}}_{K\text{ times}}.$
@@
The parameter identifiability of such networks is a complementary question, as it can only hold when this dimension is sufficiently small (similarly to tensors of subgeneric ranks). It was shown in [40] that the identifiability of deep PNNs can be characterized by the identifiability of shallow PNNs (and, thus, of tensor decompositions): a key result is that a deep ($L$-layer) polynomial NN is identifiable if and only if every every 2-layer PNN block composed by a pair of two successive layers is also identifiable.
@@ V-B1 Tensor decompositions as NNs
$f(\boldsymbol{x})$ $\,=\,\langle\mathcal{A},\phi(x_{1})\otimes\cdots\otimes\phi(x_{{d_{\rm in}}})\rangle$
$\,=\,\sum_{k_{1},\dots,k_{{d_{\rm in}}}=1}^{K}\mathcal{A}_{k_{1},\dots,k_{{d_{\rm in}}}}\prod_{i=1}^{{d_{\rm in}}}\phi_{k_{i}}(x_{i})\,$, (8)
@@ V-B3 Tensorization and approximation theory in terms of smoothness spaces
$f(i_{1},\ldots,i_{Q},v)$ $\,=\,\sum_{k=1}^{K}\zeta_{1}^{(k)}(i_{1})\dots\zeta_{Q}^{(k)}(i_{Q})\psi^{(k)}(v)$
$\,=\,\langle\mathbb{1},\zeta_{1}(i_{1})\otimes\dots\otimes\zeta_{Q}(i_{Q})\otimes\psi(v)\rangle\,$, (9)
@@ VI-2 Score function approaches
One difficulty with exploiting (10) is the need to differentiate $f(\boldsymbol{x})$. Nonetheless, we can still construct tensors to train NNs using knowledge about the input data distribution $p(\boldsymbol{x})$ by means of the so-called score function $\mathcal{S}_{m}(\boldsymbol{x})=(-1)^{m}[\nabla^{m}p(\boldsymbol{x})]/p(\boldsymbol{x})$, which is an order-$m$ tensor of size ${d_{\rm in}}\times...\times{d_{\rm in}}$ proportional the derivative of $\log p(\boldsymbol{x})$ [6]. The key idea comes from a generalization of the classical Stein’s lemma (Theorem 6 in [45]), which relates the (cross) moments between any continuously differentiable labeling function $f(\boldsymbol{x})$ and the the score function $\mathcal{S}_{m}(\boldsymbol{x})$, to the expectation of the $m$-th order derivative of $f(\boldsymbol{x})$ as
@@ VI-A2 Connecting the method of moments to optimization approaches
Tensorial approaches have been used in tandem with optimization strategies to train NNs. By analyzing the least squares loss function for learning 2-layer NNs with Gaussian inputs and no bias terms, it was shown in [48] that the optimization loss is implicitly equivalent to simultaneously decomposing a sequence of low-rank tensors, weighted by the Hermite coefficients of the activation function $\xi_{k}$, that is, $\amsmathbb{E}_{\boldsymbol{x}\sim p(\boldsymbol{x})}\{|f(\boldsymbol{x})-\hat{f}(\boldsymbol{x})|^{2}\}=\sum_{k=1}^{\infty}\xi_{k}\|\sum_{r=1}^{d_{1}}\boldsymbol{a}_{r}^{\otimes k}-\sum_{r=1}^{d_{1}}\hat{\boldsymbol{a}}_{r}^{\otimes k}\|^{2}$ for two networks $f$ and $\hat{f}$. Using this connection, the authors modified the loss function (effectively changing the activation function) to obtain more favorable optimization landscapes. The method of moments was also used to provide sufficiently accurate initializations to stochastic gradient algorithms such that convergence to a global optimum could be guaranteed [49].
@@ VI-B Using direct differentiation and flexible activation functions
$\mathcal{J}_{:,:,n}$ $=\nabla_{\boldsymbol{x}}^{(1)}f(\boldsymbol{x}_{n})$ (12)
$=\boldsymbol{A}_{2}\operatorname{diag}\big{(}\sigma^{\prime}(\boldsymbol{A}_{1}\boldsymbol{x}_{n}+\boldsymbol{b}_{1})\big{)}\boldsymbol{A}_{1}\,$, (13)
@@ VI-B2 Learning 3-layer NNs
where $\boldsymbol{t}_{n}=\boldsymbol{A}_{1}\boldsymbol{x}_{n}$ and $\boldsymbol{z}_{n}=\boldsymbol{A}_{2}\sigma_{1}(\boldsymbol{A}_{1}\boldsymbol{x}_{n})$. This is no longer an additive decomposition, but a Paratuck-2 decomposition [3]. This result was used in [23] to devise a learning algorithm with parameter identification guarantees based on a coupled factorization of the Jacobian and Hessian tensors of $f(\boldsymbol{x})$.

### v2

@@ I-1 NNs with low-rank weights
Many works represent weights in different NN architectures as low-rank tensors for their compression [8], and also for the compression of their gradient updates (such as in LoRA [9]). We will present results on the sensitivity and convergence of NN weight compression algorithms, on the generalization performance of such models, and on their implicit biases. This will be covered in Section IV.
@@ I-3 Learning with differentiation
Through the connection between NNs and tensors obtained by derivatives, the theory of low-rank decompositions helps to establish polynomial time training algorithms, generalization performance guarantees, handling trainable activation functions, and parameter identifiability for NNs. We will discuss results for learning 2- and 3-layer NNs that use this connection by leveraging different low-rank matrix and tensor formats [16, 17, 18, 19]. This will be covered in Section VI.
@@ II-D1 Generic properties
A wealth of results exist on generic properties of tensor decompositions, that is, the properties which hold for almost all tensors (as explained below) of a given size or tensor format. Generic properties can be studied using the powerful tools of algebraic geometry [24]. A key example is the generic tensor rank: all complex-valued tensors of given size $I_{1}\times\cdots\times I_{d}$ have the same CP rank $r_{\rm gen}\triangleq r_{\rm gen}(I_{1},\ldots,I_{d})$, except for a set of Lebesgue measure zero. In other words, if a complex tensor of a given size is drawn according to continuous probability distribution, then with probability one the tensor has CP rank $r_{\rm gen}$. For example, for cubic tensors $r_{\rm gen}(I,I,I)=\big\lceil\frac{I^{3}}{3I-2}\big\rceil$, and for many different tensor shapes the generic ranks are known. The generic ranks also give a bound $r_{\rm max}\leq 2r_{\rm gen}$ on maximal rank $r_{\rm max}$ of a given tensor size, which is otherwise difficult to study and unknown for most cases. For real-valued tensors, several typical ranks can occur (corresponding to sets of positive Lebesque measure), but the smallest typical rank is equal to $r_{\rm gen}$, see [2, 25, 24] for an overview.
@@ II-D3 Other factorizations
Despite being widely used to explain the behavior of NNs, the CPD has limited flexibility, which is one reason many results only address the 2-layer case (see, e.g., [16, 29]). Different generalizations of the CPD exist: symmetric, partially symmetric (where symmetry is imposed on factors of the rank-one components), block-term decomposition (BTD, decomposition of a tensor as a sum of low-rank tensors); and the so-called tubal tensor rank decomposition, or t-SVD (often used for multichannel/3D data) which is a special case of the BTD with a fixed factor matrix [30]. The above decompositions are additive, and most of them fall under the umbrella of X-rank decompositions [25, 24]. Many generic properties of such decompositions (e.g., identifiability) can be studied using the tools from algebraic geometry, similarly to the case of CPD. Besides additive decompositions, there are other generalizations of the CPD such as the Paratuck-2 [3], which enjoys uniqueness guarantees and can be useful in the analysis of deep learning models.
@@ III Neural networks: basic setup
$\boldsymbol{h}_{\ell}\,=\,f_{\ell}(\boldsymbol{h}_{\ell-1})\,\triangleq\,\sigma_{\ell}(\boldsymbol{A}_{\ell}\boldsymbol{h}_{\ell-1}+\boldsymbol{b}_{\ell})\,,\qquad\text{(hidden layers)}$
@@
CP $\bullet\bullet$
NP-hard in general but well studied Known generic ranks (linked to expressivity of NNs); NN learning with derivatives and flexible activations
Symmetric CP $\bullet$
Good polynomial time algorithms for some rank values Stable under perturbations; key tool for learning NNs with derivatives/method of moments
Tucker $\bullet\bullet\bullet$
Efficient algorithms based on linear algebra Widely used for compression of NN weights
TT and HT $\bullet\bullet\bullet$
Efficient algorithms based on linear algebra Known generic ranks (linked to expressivity of NNs); well-adapted to work in high-dimensions; directly linked to (the study of) sum-product architectures
Paratuck $\bullet\bullet$
Hard to compute Linked to deep (>2>2 layers) NNs by derivatives
@@ IV-A Compression of weights
Note that tensorization has been used not only to compress NNs but also to increase their expressiveness, as a means of capturing higher order interactions between inputs/hidden variables. This was used in tensor attention modules in transformers and in second-order RNNs [11], which were proven to have increased expressiveness over their original counterparts. However, the most common use of low-rank representations is still to reduce not only storage costs but also training and inference time. This is because in many cases the computations can be performed using the decomposed format itself, without need to reconstruct the full weight tensor (which would not decrease the computation time during inference).
The availability of different low-rank formats introduces the question of what is the best loss function or tensor rank for compressing a given NN architecture. Relevant work in this direction include using stable loss functions to avoid degenerate components when compressing CNN layers (see [5]), and [32], which uses the Riemannian geometry of the Tucker decomposition to train NNs in compressed format with improved convergence performance, besides proposing an adaptive rank selection strategy.
@@ V-A2 Polynomial neural networks
$f(\boldsymbol{x})=\boldsymbol{A}_{2}\sigma_{1}(\boldsymbol{A}_{1}\boldsymbol{x})\,,\iff\mathcal{F}\emph{ has a CPD with factors }\boldsymbol{A}_{2},\underbrace{\boldsymbol{A}_{1}^{\top},\ldots,\boldsymbol{A}^{\top}_{1}}_{K\text{ times}}.$
@@
The parameter identifiability of such networks is a complementary question, as it can only hold when this dimension is sufficiently small (similarly to tensors of subgeneric ranks). It was shown in [14] that the identifiability of deep PNNs is intimately linked to the identifiability of shallow PNNs (and, thus, of tensor decompositions): a key result is that a deep ($L$-layer) polynomial NN is identifiable if every every 2-layer PNN subnetwork composed by a pair of two successive layers is also identifiable.
@@ V-B1 Tensor decompositions as NNs
$f(\boldsymbol{x})$ $\,=\,\langle\mathcal{A},\phi(x_{1})\otimes\cdots\otimes\phi(x_{{d_{\rm in}}})\rangle\,=\,\sum_{k_{1},\dots,k_{{d_{\rm in}}}=1}^{K}\mathcal{A}_{k_{1},\dots,k_{{d_{\rm in}}}}\prod_{i=1}^{{d_{\rm in}}}\phi_{k_{i}}(x_{i})\,$, (7)
@@ V-B3 Tensorization and approximation theory in terms of smoothness spaces
$f(i_{1},\ldots,i_{Q},v)$ $\,=\,\sum_{k=1}^{K}\zeta_{1}^{(k)}(i_{1})\dots\zeta_{Q}^{(k)}(i_{Q})\psi^{(k)}(v)\,=\,\langle\mathbb{1},\zeta_{1}(i_{1})\otimes\dots\otimes\zeta_{Q}(i_{Q})\otimes\psi(v)\rangle\,$, (8)
@@ VI-2 Score function approaches
One difficulty with exploiting (9) is the need to differentiate $f(\boldsymbol{x})$. Nonetheless, we can still construct tensors to train NNs using knowledge about the input data distribution $p(\boldsymbol{x})$ by means of the so-called score function $\mathcal{S}_{m}(\boldsymbol{x})=(-1)^{m}[\nabla^{m}p(\boldsymbol{x})]/p(\boldsymbol{x})$, which is an order-$m$ tensor of size ${d_{\rm in}}\times...\times{d_{\rm in}}$ proportional the derivative of $\log p(\boldsymbol{x})$ [6, 16]. The key idea comes from a generalization of the classical Stein’s lemma, which relates the (cross) moments between any continuously differentiable labeling function $f(\boldsymbol{x})$ and the the score function $\mathcal{S}_{m}(\boldsymbol{x})$, to the expectation of the $m$-th order derivative of $f(\boldsymbol{x})$ as
@@ VI-A2 Connecting the method of moments to optimization approaches
Tensorial approaches have been used in tandem with optimization strategies to train NNs. By analyzing the least squares loss function for learning 2-layer NNs with Gaussian inputs and no bias terms, it was shown in [43] that the optimization loss is implicitly equivalent to simultaneously decomposing a sequence of low-rank tensors, weighted by the Hermite coefficients of the activation function $\xi_{k}$, that is, $\amsmathbb{E}_{\boldsymbol{x}\sim p(\boldsymbol{x})}\{|f(\boldsymbol{x})-\hat{f}(\boldsymbol{x})|^{2}\}=\sum_{k=1}^{\infty}\xi_{k}\|\sum_{r=1}^{d_{1}}\boldsymbol{a}_{r}^{\otimes k}-\sum_{r=1}^{d_{1}}\hat{\boldsymbol{a}}_{r}^{\otimes k}\|^{2}$ for two networks $f$ and $\hat{f}$. Using this connection, the authors modified the loss function (effectively changing the activation function) to obtain more favorable optimization landscapes. The method of moments can be also used to provide sufficiently accurate initializations to stochastic gradient algorithms such that convergence to a global optimum could be guaranteed.
@@ VI-B Using direct differentiation and flexible activation functions
$\mathcal{J}_{:,:,n}\,\triangleq\,\nabla_{\boldsymbol{x}}^{(1)}f(\boldsymbol{x}_{n})\,=\,\boldsymbol{A}_{2}\operatorname{diag}\big(\sigma^{\prime}(\boldsymbol{A}_{1}\boldsymbol{x}_{n}+\boldsymbol{b}_{1})\big)\boldsymbol{A}_{1}\,$,
@@ VI-B2 Learning 3-layer NNs
where $\boldsymbol{t}_{n}=\boldsymbol{A}_{1}\boldsymbol{x}_{n}$ and $\boldsymbol{z}_{n}=\boldsymbol{A}_{2}\sigma_{1}(\boldsymbol{A}_{1}\boldsymbol{x}_{n})$. This is no longer an additive decomposition, but a Paratuck-2 decomposition [3]. This result was used in [44] to devise a learning algorithm based on the factorization of the Jacobian of $f(\boldsymbol{x})$.
@@ VII-3 Reinforcement learning
Acknowledgment
This work was supported in part by the French National Research Agency (ANR) under grants ANR- 23-CE23-0024, ANR-23-CE94-0001, by the PEPR project CAUSALI-T-AI, and by the National Science Foundation, under grant NSF 2316420.